OpenStudy (acespeedfighter):

what is the value of 1/3 x^2 + 5.2y when x = 3 and y = 2?

OpenStudy (freckles):

replace x with 3 and replace y with 2

OpenStudy (freckles):

then apply order of operations

OpenStudy (acespeedfighter):

wait how do i do that?

OpenStudy (acespeedfighter):

@_citrus_

OpenStudy (freckles):

\[\frac{1}{3} \color{red}{x}^2+5.2 \color{blue}{y}\]

OpenStudy (freckles):

the x is the one in red the y is the one in blue just coloring them so you can see the x and y clearly

OpenStudy (freckles):

so the red x will be replaced with 3 and the blue y will be replace with 2

OpenStudy (acespeedfighter):

6 1/3 + 5.22?

OpenStudy (freckles):

1/3 x^2 means 1/3*x^2 and 5.2y means 5.2*y

OpenStudy (freckles):

\[\frac{1}{3}(x^2)+5.2(y) \\ \frac{1}{3}(3)^2+5.2(2)\]

OpenStudy (acespeedfighter):

OH

OpenStudy (freckles):

take care of exponents then the multiplication then do the addition

OpenStudy (acespeedfighter):

so 1/3 x 6

OpenStudy (freckles):

3^2 means 3*3 not 3*2

OpenStudy (acespeedfighter):

and then 5.2 x 2

OpenStudy (acespeedfighter):

XD i meant 1/3 * 9 right?

OpenStudy (freckles):

yep

OpenStudy (freckles):

1/3*9=? and 5.2*2=?

OpenStudy (acespeedfighter):

then add those 2 numver?

OpenStudy (freckles):

yep add those results and you be done

OpenStudy (acespeedfighter):

13.4?

OpenStudy (freckles):

looks dandy

OpenStudy (acespeedfighter):

So am i right :D

OpenStudy (freckles):

yep

OpenStudy (acespeedfighter):

:D Thx @pooja195

pooja195 (pooja195):

You mean @freckles

OpenStudy (acespeedfighter):

no i mean @freckles

OpenStudy (acespeedfighter):

you know what i mean :D

OpenStudy (freckles):

lol

OpenStudy (freckles):

np